# SonicVisionLM: Playing Sound with Vision Language Models

### v2

@@ 4 Experiments
4.1 Experiment Settings.
Figure 4: Conditional Generation Task Qualitative Results. The red dashed boxes are the conditional audio inputs and generated results for the CondFoleyGen task, and the blue dashed boxes are the conditional text inputs and generated results corresponding to SonicVisionLM.
Dataset. Considering that Tango is a large model that incorporates various types of audio data, including speech, sound effects, and music, we have developed a high-quality single sound effects dataset called CondPromptBank. This dataset is specifically designed for training a time-controllable adapter, which allows us to align the generated sound with our task requirements while ensuring precise timing control.
CondPromptBank consists of 10,276 individual data entries. Each entry includes a high-quality sound effect, its corresponding sound effect title, and the starting and ending timestamps for when the sound is emitted. The duration of these sound effects is typically 10 seconds or shorter. All the data in CondPromptBank is sourced from freely available sound effect libraries and websites. During collection, we focused on 23 common categories of sound effects and manually filtered out low-quality data with high levels of noise and mixed sources.
To provide comprehensive textual descriptions, we further annotated the sound effect text labels with fine-grained information based on sound characteristics. Each sound effect text label now includes details about the single audio source and the associated action. We believe that this division into single audio sources is crucial because videos, as compared to text, provide a more concrete expression of content, demanding a greater accuracy in sound representation.
Implementation details. To train our complete model, we first train the adapter on CondPromtBank, then train the timestamps detection net using a large video-audio-time dataset. We trained the adapter for approximately 200 epochs with a batch size of 32 and a learning rate of $3.0\times 10^{-5}$ using Adam [16]. Our modal has two versions: small and full, the difference between them is the versions of pre-trained parameters used for Tango. We trained the timestamps detection net for 70 epochs with a batch size of 24 and a learning rate of $1.0\times 10^{-5}$. The training of the adapter takes approximately 5 days and the training of the the timestamps detection network requires approximately 1 day using 1 NVIDIA A6000 GPU.
@@
Table 2: Unconditional Sound Generation Quantitative Results. IS assesses the quality and diversity of generated samples, FID measures distribution-level similarity, and MKL measures paired sample-level similarity. $*{}^{*}$ ”denotes the data sourced from the official code and is based on experiments conducted with our local configurations. underline denotes the worst performance. boldface denotes the best performance. IoU metric is measured in percentage.
@@ 4.2 Conditional Generation Task Results
Evaluation Metrics. For the conditional generation task, we use the following five objective metrics to evaluate the performance of the model: CLAP-top1, Onset Acc[8], Onset AP[8], Time Acc, and IoU.
Qualitative Results. As seen in Tab.1, SonicVisionLM-full demonstrates superior performance over CondFoleyGen across all metrics. It got a CLAP-top1 score of 46.7%, 30% higher than any versions of the CondFoleyGen model This result suggests that it’s better at matching sounds to text prompts. In audio-visual sync, our model outperforms CondFoleyGen significantly, achieving an Onset AP score that is 11% higher and an IoU score that is 16% higher. These results underscore our model’s enhanced ability to generate sounds that are more accurately synchronized with the input video in terms of timing. Our model outperforms CondFoleyGen in onset and timing accuracy, showing improvements of 6% and 8.3%, respectively. This underscores its precision in detecting sound events and its effectiveness in aligning visual and audio elements.
We think CondFoleyGen relies on the audio-visual synchronization module[14] to improve time accuracy, which needs a large number of samples to re-ranking the sounds. In contrast, our model gets higher synchronization during the generation process, with fewer samples.
Quantitative Results. As shown in Fig.4 (row 7,8,9), we compare the timestamped positional distance to the target sound. Our results closely match the target sound, demonstrating a high degree of accuracy. However, CondFoleyGen often produces results with the wrong number of sounds and with a large difference in the positional distance This indicates that our model has a higher visual-audio synchronization than CondFoleyGen. As shown in Fig.4 (column 3,7), we compare the sound shape to the conditional sound. Even though our model is not trained on the Greatest Hits dataset, our results are still very similar to the conditional sound. As shown in Fig.4 (column 1,7), The mel-spectrograms of the CondFoleyGen results do not produce sound that are similar to, and sometimes blurred.
@@ 4.3 Unconditional Generation
To evaluate the task of unconditional sound generation, considering that our model has not been trained on audio-visual datasets, we have chosen to perform quantitative evaluation and qualitative evaluation on two datasets: Greatest Hits[25] and CountixAV[40], which are zero-shot tasks for all models. We use two state-of-the-art condition V2A models: SpecVQGAN [13] and DIFF-FOLEY[24] as baselines.
Evaluation Metrics. For objective evaluation, we have employed three metrics as [13, 24]: Inception Score (IS)[30], Frechet Distance (FID)[10], and Mean KL Divergence (MKL)[13].
For subjective evaluation, As the [8], we conduct user evaluations for the three critical components: overall audio quality (OVL), alignment with the input video (REL), and time synchronization (Time-sync).
@@
Table 3: Subjective Results. Following [8], we invited the 300 evaluators to judge 30 randomly selected audio samples from three perspectives: OVL, REL, and Time-sync. Ratings were scored on a scale from 1 to 100. All evaluators were proficient in English. The scoring scale is 1-100 and results are averaged out. $*{}^{*}$ ”denotes the data sourced from the official code and is based on experiments conducted with our local configurations. underline denotes the worst performance. boldface denotes the best performance.
Quantitative Results. As shown in Tab.2, our model notably surpasses DIFF-FOLEY and SpecVQGAN on the Greatest Hits dataset in terms of MKL and FID metrics, with respective improvements of 1.76, 6.4, and 2.13, 57.5. This signifies our model’s superior ability to generate sounds that align closely with the ground truth. On the CountixAV dataset, SonicVisionLM-small is only marginally behind the best result in the IS metric by 0.2, suggesting a commendable sound quality produced by our model. Despite our model’s slightly lower performance in the MKL metric compared to the baseline, we view this not as a drawback. Unlike the Greatest Hits dataset, which was recorded in high quality with specialized recording equipment, the CountixAV dataset is sourced from diverse YouTube videos, and often includes sounds marred by low-quality background noise or non-homogenous sound sources. The MKL and FID metrics, focusing on similarity to ground truth, which could explain why SpecVQGAN, despite primarily generating noisy sounds, excels in the MKL metric. The IoU metrics show that our model makes sound at the right times and keeps quiet at other times. To further argue our point, we conducted a subjective assessment of three aspects of OVL, REL, and Time-sync. We found that our model performed significantly better than the baseline models in all metrics, as shown in Table.3
Qualitative Results. To compare the time points more visually, we denoised the CountixAV dataset. As we can see from Fig.5, our generated sounds are more clear than the baseline models. On top of that, our model can capture finer-grained visual transformations, capturing the signal despite sudden changes in the frequency of movement and producing the correct sound. in contrast, SpecVQGAN results have only noise throughout the sound. DIFF-FOLEY sometimes can generate sounds without noise, but it is not aligned with the visual-sound synchronization.
@@
Table 4: Time Condition Ablation $*{}^{*}$ ”denotes the data sourced from the official code and is based on experiments conducted with our local configurations. underline denotes the worst performance. boldface denotes the best performance. IoU metric is measured in percentage.
Figure 6: Conditioning-scale Ablation Study. In the chart, the higher the indicators are, the better they are. The brightness of the color also represents a good or bad performance.
@@ 4.4 Ablation Study
Time condition. we compared our model with the tango[9]. Evaluation Metrics. In the ablation study, We employed objective metrics from Tango[9], FAD [15] and Distance (FD)[23] measure the distribution difference between generated audio and real audio without the need for any reference audio samples. KL divergence [37, 18] assesses the similarity between them based on the broad concepts present in the original and the generated audio signals.
Result. As shown in Tab.4, both in the small and full versions, our model outperforms tango in terms of IoU, which validates the effectiveness of the time-controllable adapter.
Conditioning-scale. As seen in the Fig.6. In the range from 0.6 to 2.0 on the conditioning scale, the sound quality improves as it lines up better with the IoU. But at 3.0, control goes down a lot. The best mix of good sound quality and steadiness is found at 2.0.
@@ 4.5 Multi-soundtracks generation
As shown in Fig.7, SonicVisionLM works like a sound designer in traditional video post-production. First, it carefully looks at the visual information to find the key sounds needed. For example, it picks out the hissing of cooking meat and the clinking of pots. Then, if the story needs it, the sound designer adds off-screen sounds. This might include extra kitchen sounds like distant chopping, simmering pots, or the buzz of kitchen machines, making it feel more real. Lastly, SonicVisionLM puts these audio tracks in the right place on the video timeline. This makes sure the sounds match up perfectly with what’s happening in the video, creating a smooth audio-visual experience. More detailed results can be found in the Supplementary Materials.
Figure 7: Multi-soundtracks generation example. In the video, the blue and green masks show different soundtracks. You can notice that the timing of on-screen sound in the video (like actions happening) matches when they happen in the frame. But the sounds that come from off-screen (not seen in the video) don’t have precise timing.
@@ 5 Conclusion
In this paper, we propose SonicVisionLM, which utilizes the capabilities of powerful vision language models (VLMs). When provided with a silent video, SonicVisionLM first identifies events within the video using a VLM to suggest possible sounds that match the video content. SonicVisionLM demonstrates outstanding performance in both conditional and unconditional generation tasks. Additionally, we tested its efficacy in the post-production process, focusing on automatic recognition of on-screen sounds and personalized editing for off-screen sounds. Extensive experiments show superior performance of our method.
Limitations. While SonicVisionLM has achieved satisfactory results, refinements in the visual understanding and timestamp detection parts are still required. When it comes to various multimedia contexts, expanding the diversity and range of audio generation control are still the key demands of applications. Thus, the model should enrich the audio-visual experience and broaden applicability in future attempts.
@@
6 Timestamp Detection Module Precision experiment
Our timestamp detection model shows promising performance, with an Average Precision(AP) of 0.80 and an accuracy of 0.72 on the Greatest Hits test set, and even higher results on the validation set with an AP of 0.92 and accuracy of 0.82. However, its performance on the CountixAV test set is comparatively lower, achieving an AP of 0.52 and accuracy of 0.52. This discrepancy likely stems from the complexity of sound sources in our training dataset, leading to potential inaccuracies in ground truth (GT). Such labeling challenges can adversely affect recognition accuracy, particularly in scenarios involving non-static footage.
7 Subjective Results Details
We provide the screenshot of the main evaluation page that the participant will see during the test in Fig.8.
@@ 7 Subjective Results Details
Figure 9: Screenshot of the conditional generation task section in the demo video.
8 Additional Results
Figure 10: Screenshot of the unconditional generation task section in the demo video.
@@ 8 Additional Results
Conditional Generation Results As shown in Fig.9, the left column represents the target video, the middle column showcases the control conditions for CondFoleyGen and our model, while the right column displays the generated results. We provide 6 examples for the previously mentioned conditional generation task, and the corresponding audio outcomes can be observed in the demo video.
Unconditional Generation Results As shown in Fig.10, we compare a comparison of the results generated by GT, SpecVQGAN, DIFF-FOLEY, and our model. We provide 6 examples from CountixAV, and the audio samples are available in the demo video.
Multi-soundtracks Generation Results As shown in Fig.11, we segment the video into different shots firstly. Shots with clear actions are directly processed by the VLMs to obtain corresponding sound effect descriptions and their spatial positioning within the video. These descriptions are used to generate on-screen sound. Other shots accept user editing and are then fed into the LDM to produce off-screen sounds.We provide a simple and a complex case respectively in the demo video.

### v3

@@ 4 Experiments
4.1 Experiment Settings
Figure 4: Conditional Generation Task Qualitative Results. The red dashed boxes are the conditional audio inputs and generated results for CondFoleyGen, and the blue dashed boxes are the conditional text inputs and results corresponding to SonicVisionLM.
Dataset. Since Tango’s training datasets include varied audio types like speech, sound effects, and music, it tends to produce mixed audio outputs. Our task requires distinct handling of these audio types, ensuring each sound event is separate. To meet our specific needs, we developed CondPromptBank, a high-quality dataset of single sound effects crafted for training time-controlled adapters. This dataset consists of 10,276 individual data entries, each with a sound effect, title, and start/end timestamps. Each sound is typically 10 seconds or shorter, sourced from freely available sound effect libraries and websites. During the collection, we focused on 23 common categories of sound effects and manually filtered out low-quality data with noise and mixed sources. The category distribution shows Tab. 1. To enhance the textual descriptions with precise details, we further annotated the sound effect text labels with fine-grained information based on sound characteristics. Now, each label not only identifies the audio source but also describes the associated actions in detail. We believe the division into single audio sources is crucial, as videos provide a more concrete expression of content than text, thus requiring more precise sound representation.
Implementation details. To train our complete model, we first train the adapter on CondPromtBank, then train the timestamps detection net on Greatest Hits [26] and Countix-AV dataset [41]. We trained the adapter for approximately 200 epochs with a batch size 32 and a learning rate of $3.0\times 10^{-5}$ using Adam [17]. Our model has two versions: small and full. The difference between them is the version of pre-trained parameters used for Tango. We trained the timestamps detection net for 70 epochs with a batch size of 24 and a learning rate of $1.0\times 10^{-5}$. The training of the adapter takes approximately five days, and the training of the timestamps detection network requires approximately one day using one NVIDIA A6000 GPU.
@@ 4.1 Experiment Settings
Table 3: Unconditional Sound Generation Quantitative Results. IS assesses the quality and diversity of generated samples, FID measures distribution-level similarity, and MKL measures paired sample-level similarity. “ ”denotes the data sourced from the official code and is based on experiments conducted with our local configurations. “ ”denotes that data is obtained from our adjusted official code. Underline denotes the worst performance. Boldface denotes the best performance. IoU metric is measured in percentage.
@@ 4.2 Conditional Generation Task Results
Evaluation Metrics. For the conditional generation task, we use the following five objective metrics to evaluate the performance of the model: CLAP-top, Onset Acc [9], Onset AP [9], Time Acc, and IoU.
Quantitative Results. SonicVisionLM-full demonstrates superior performance over CondFoleyGen across all metrics, as seen in Tab. 2. It achieves leading scores of 36.8% and 42.8% on CLAP-top versions, exceeding any CondFoleyGen model by more than 20%. This result suggests that it is better at matching sounds to text prompts. Our model significantly surpasses CondFoleyGen in audio-visual synchronization, evidenced by an 11% improvement in Onset AP and a 16% rise in IoU. Moreover, we see a 6% enhancement in onset accuracy and an 8.3% increase in timing precision. These advancements illustrate our model’s outstanding accuracy in sound event detection and its effectiveness in synchronizing generated sounds with the input video. We think CondFoleyGen relies on the audio-visual synchronization module [15] to improve time accuracy, requiring many samples to re-ranking the sounds. In contrast, our model gets higher synchronization during the generation process with fewer samples.
Qualitative Results. As shown in Fig. 4 (row 4, 6, 7), we compare the timestamped positional distance to the target sound. Our results closely match the target sound, demonstrating a high degree of accuracy. However, CondFoleyGen often produces results with the wrong number of sounds and with a large difference in the positional distance. This indicates that our model has a higher visual-audio synchronization than CondFoleyGen. As shown in Fig. 4 (column 3, 7), we compare the sound shape to the conditional sound. Even though our model has not learned any timbre on the Greatest Hits dataset, our results are still very similar to the conditional sound. As shown in Fig. 4 (column 1, 7), The mel-spectrograms of the CondFoleyGen results do not produce sounds that are similar to, and sometimes blurred.
@@ 4.3 Unconditional Generation
To evaluate the task of unconditional sound generation, considering that our LDM has not been trained on audio-visual datasets, we have chosen to perform quantitative evaluation and qualitative evaluation on two datasets: Greatest Hits [26] and CountixAV [41], which are zero-shot tasks for all models. We use two state-of-the-art V2A models as baselines: SpecVQGAN [14] and DIFF-FOLEY [25].
Evaluation Metrics. For objective evaluation, we have employed three metrics as [14, 25]: Inception Score (IS) [31], Frechet Distance (FID) [11], and Mean KL Divergence (MKL) [14]. For subjective evaluation, as the [9], we conduct user evaluations for the three critical components: overall audio quality (OVL), alignment with the input video (REL), and time synchronization (Time-sync).
@@
Table 4: Subjective Results. Following [9], we invited the 300 English-proficient evaluators to rate 30 randomly selected audio samples from three perspectives: OVL, REL, and Time-sync. Scores were averaged on a 1-100 scale. “” denotes the data sourced from the official code and is based on experiments conducted with our local configurations. “” denotes data obtained from our adjusted official code. Underline denotes the worst performance. Boldface denotes the best performance.
Quantitative Results. In our experiments, the choice of vocoder significantly influenced the FID metric. We upgraded from DIFF-FOLEY’s weak Griffin-Lim vocoder to the superior MelGAN to ensure fairness. Despite our model using 64 Mel filters compared to MelGAN’s 80, our performance on the Greatest Hits dataset excels in MKL, IS, and IoU metrics, as shown in Tab. 3. This underscores our model’s precise time control and high-quality sound generation, demonstrating its exceptional capability to produce sounds that align closely with the ground truth. In the CountixAV dataset, SonicVisionLM-small outperforms baselines by nearly 10 points in the IS metric, highlighting its exceptional sound quality. Despite slightly lower MKL and FID scores compared to baselines, we do not view this as a drawback. Unlike the Greatest Hits dataset, which was recorded in high quality with specialized recording equipment, the CountixAV dataset is sourced from diverse YouTube videos. It often includes sounds marred by low-quality background noise or mixed sound sources, especially human vocals. MKL and FID metrics emphasize similarity to the ground truth. Since DIFF-FOLEY was trained on similar audio-visual data, its tendency to mix multiple sounds in response to complex visual information explains its advantage in these metrics. Our model, targeting sound events from specific actions and filtering out irrelevant auditory information, diverges from the ground truth, a deliberate choice to enhance sound event relevance. The IoU metrics show that our model makes sound at the correct times and remains silent otherwise. To further argue our point, we conducted a subjective assessment of three aspects of OVL, REL, and Time-sync. As shown in Tab. 4, our model performed significantly better than the baseline models in all metrics.
Qualitative Results. For clearer time point comparison, we denoised the data from the CountixAV. Fig. 5 reveals our generated sounds are more precise than the baseline models. Furthermore, our model can capture finer-grained visual transformations, capturing the signal despite sudden changes in the frequency of movement and producing the correct sound. In contrast, SpecVQGAN results are full of noise. DIFF-FOLEY can sometimes generate clear sounds, but it is not aligned with the visual-sound synchronization.
@@
Table 5: Time Condition Ablation Results. ”denotes the data sourced from the official code and is based on experiments conducted with our local configurations. Underline denotes the worst performance. Boldface denotes the best performance. IoU metric is measured in percentage.
Figure 6: Conditioning-scale Ablation Study. In the chart, the higher the indicators are, the better they are. The brightness of the colour also represents a good or bad performance.
Figure 7: Multi-soundtracks generation example. In the video, the blue and green masks show different soundtracks. The timing of on-screen sounds (corresponding to visible actions) aligns accurately with their occurrence in the frames. Conversely, off-screen sounds (not visible in the video) lack precise timing.
@@ 4.4 Ablation Study
Time condition. Evaluation Metrics. We employed objective metrics from Tango [10], such as the Frechet Audio Distance (FAD) [16] and Frechet Distance(FD) [24] to measure the distribution difference between generated audio and ground truth without the need for any reference audio samples. KL divergence [38, 19] assesses their similarity based on the broad concepts present in the original and the generated audio signals.
Result. we compared our model with the Tango [10]. As shown in Tab. 5, both in the small and full versions, our model outperforms Tango in terms of IoU, which validates the effectiveness of the time-controllable adapter.
Conditioning-scale. As seen in the Fig. 6. The sound quality improves in the range from 0.6 to 2.0 on the conditioning scale as it lines up better with the IoU. But at 3.0, control goes down a lot. The best mix of good sound quality and steadiness is found at a conditioning scale equal to 2.0.
@@ 4.5 Multi-soundtracks generation
As shown in Fig. 7, SonicVisionLM works like a sound designer in traditional video post-production. First, it looks at the visual information to find the key sounds needed. For example, it picks out the hissing of cooking meat and the clinking of pots. Then, if the story needs it, the sound designer adds off-screen sounds. This might include extra kitchen sounds like distant chopping, simmering pots, or the buzz of kitchen machines, making it feel more natural. Lastly, SonicVisionLM puts these audio tracks in the right place on the video timeline. This ensures the sounds match perfectly with what is happening in the video, creating a smooth audio-visual experience. More detailed results can be found in the Supplementary Materials.
@@ 5 Conclusion
In this paper, we propose SonicVisionLM, which utilizes the capabilities of powerful vision-language models (VLMs). When provided with a silent video, SonicVisionLM first identifies events within the video using a VLM to suggest possible sounds that match the video content. SonicVisionLM demonstrates outstanding performance in both conditional and unconditional generation tasks. Further, we tested its efficacy in post-production, focusing on automatic recognition of on-screen sounds and personalized editing for off-screen sounds. Extensive experiments show the superior performance of our method.
Limitations. While SonicVisionLM has achieved adequate results, refinements in the visual understanding and timestamp detection parts are still required. Regarding various multimedia contexts, expanding the diversity and range of audio generation control are still the key demands of applications. Thus, the model should enrich the audio-visual experience and broaden applicability in future attempts.
@@
A.1 Subjective Results Details
We provide the screenshot of the main evaluation page the participant will see during the test in Fig.8.
A.2 Timestamp Detection Module Precision Experiment
Our timestamp detection model shows promising performance, with an Average Precision(AP) of 0.80 and an accuracy of 0.72 on the Greatest Hits test set, and even higher results on the validation set with an AP of 0.92 and accuracy of 0.82. However, its performance on the CountixAV test set is comparatively lower, achieving an AP of 0.52 and an accuracy of 0.52. This discrepancy likely stems from the complexity of sound sources in our training dataset, leading to potential inaccuracies in ground truth. Such labelling challenges can adversely affect recognition accuracy, particularly in scenarios involving non-static footage.
Figure 9: screenshot of the conditional sound generation task section in the demo video.
A.3 Additional Results
Figure 10: screenshot of the unconditional sound generation task section in the demo video.
@@ A.3 Additional Results
Conditional Generation Results. As shown in Fig.9, the left column represents the target video, the middle column showcases the control conditions for CondFoleyGen and our model, and the right column displays the generated results. We provide 6 examples for the previously mentioned conditional generation task, and the corresponding audio outcomes can be observed in the demo video.
Unconditional Generation Results. As shown in Fig.10, we compare a comparison of the results generated by GT, SpecVQGAN, DIFF-FOLEY, and our model. We provide 6 examples from CountixAV, and the audio samples are available in the demo video.
Multi-soundtracks Generation Results. As shown in Fig.11, we segment the video into different shots first. Shots with clear actions are directly processed by the VLMs to obtain corresponding sound effect descriptions and their spatial positioning within the video. These descriptions are used to generate on-screen sound. Other shots accept user editing and are then fed into the LDM to produce off-screen sounds. We provide a simple and a complex case respectively in the demo video.